OpenStudy (anonymous):

find exact value of cot 23pi/12

OpenStudy (anonymous):

\[\cot (23\pi/12)\]

OpenStudy (anonymous):

23.26 ?

OpenStudy (asnaseer):

HINT: \(\frac{23\pi}{12}=\frac{24\pi}{12}-\frac{\pi}{12}=2\pi-\frac{\pi}{12}\)